# Distance-Domain Degrees of Freedom in Near-Field Region

### v2

@@ IV-C1 Comparison with Angular-Domain DoF Expressions
In many P2P MIMO setups where the Rx aperture mainly spans the angular extent of the Tx aperture, the angular-domain DoF expressions often take the form [14, 16, 25, 12, 13, 27, 17, 21, 22, 24, 20]:
@@
where $A_{t}$ and $A_{r}$ are the physical areas of Tx and Rx arrays, $D$ is the Tx–Rx separation (treated as a single representative distance), $n$ depends on aperture dimensionality, and $\alpha$ is coefficients determined via orientation of Tx and Rx arrays. More elaborate variants appear in [19]. Our distance-domain result in (50) differs in several key ways:
@@ IV-C2 Comparison with Distance-Domain DoF in the Literature
Let us consider [28], which analyzes two perpendicular ULAs—a geometry closest to our setting. If the Tx array center lies on the Rx-line direction, their closed-form DoF can be written as [28, Eqs. (39),(33)]
@@
Figure 6: Distance-domain DoF vs length of Rx array obtained via approaches: proposed analytical DoF, [28]. and simulation using SVD (with threshold $\epsilon=0.1\sigma_{\max}$). Tx array is ULA with length of $100\lambda$, and $r_{\min}=100\lambda$, while the length of Rx array ($r_{\max}-r_{\min}$) ranges between $10\lambda,2000\lambda]$.
@@
in agreement with simulations (Fig. 6). Thus, when users are spread over a wide range of distances, our proposed DoF expression in (50) captures the correct bounded scaling, while DoF expression in (41) from [28] overestimates growth due to its implicit narrow distance-span assumption.
Remark: Unlike many P2P–MIMO DoF’s formulas that implicitly require $r_{\min}\!\approx\!r_{\max}$, our distance-domain DoF expression in (50) remains well-posed and can accommodate the large distance spread ($r_{\max}\gg r_{\min}$). This allows our distance-domain DoF expression to capture the MU-MIMO channel where user’s distance separation can be arbitrarily large.
@@ IV-C3 Non-Uniformity of the Eigenvalue Spectrum
A common (often implicit) assumption in DoF estimates is a nearly flat plateau of dominant eigenvalues followed by rapid decay. This is reasonable when the channel kernel is (approximately) convolutional and the Fourier transform of its kernel is close to an in-band indicator over a passband $\Omega$ (flat in-band) [33]:
@@
Implication: Effective-DoF method such as [34, 22, 19, 35]:
@@ V-B Approximation Accuracy for Distance-Domain DoF in the Non-Broadside Case
  • Exact phase & exact amplitude: further incorporates polarization and elementwise distance variation in amplitude [36].
@@ VI-B1 Distance-domain DoF under physical length constraint
For this special case, one can see that $\frac{(a+b)}{2}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)$ is approximately the angle subtended by the Rx array seen from the center of the linear Tx sub-array $\mathcal{P}^{(1)}$. Hence, the distance-domain DoF in this special case is analogous to the angular-domain DoF derived for the linear array, as discussed in [37].
As shown by (60), shows that for a fixed sub-array length $L$, increasing the extreme edges (i.e., increasing both $a$ and $b$) expands the distance-domain DoF. Intuitively, increasing the extreme edges expands the angle seen from the center of the sub-array, hence increasing the DoF [37]. This is illustrated in Fig. 14, which suggests that a modular array can exploit additional spatial resources in the distance domain compared to a single-piece array with the same overall physical length.
@@ VI-B2 Spatial DoF under extreme edge constraint
Figure 15: Eigenvalue distributions of LoS channel where the Tx array is either the linear array or the symmetric modular linear array (All inner gaps of the modular arrays have the same size, and the Rx array satisfies $r_{\min}=200\lambda$ and $r_{\max}=2000\lambda$).
@@ VII Conclusion
In this paper, we investigated the distance-domain DoF by considering a MIMO channel where the source (represents base station) equips with an Tx array with arbitrary shape and the observation region (represents users) is Rx linear array whose elements are aligned at the same angle but lie at different distances from the Tx array. In an ideal scenario where the Tx array is single-piece and the Rx array is on the broadside of the Tx array, we consider both the arrays having continuous aperture with infinite elements and infinitesimal spacing and model the channel as an integral operator with a non-convolution kernel. We further transform it into the integral operator with a Hermitian convolution kernel, which enables eigenvalue analysis via the Fourier transform. This formulation allowed us to derive a closed-form expression for the distance-domain DoF, demonstrating that the distance-domain DoF is mainly determined by the extreme boundaries of the Tx and Rx arrays rather than its detailed shape.

### v3

@@ IV-C1 Comparison with Angular-Domain DoF Expressions
In many P2P MIMO setups where the Rx aperture mainly spans the angular extent of the Tx aperture, the angular-domain DoF expressions often take the form [torkildson2011indoor, pizzo2022landau, miller2000communicating, xie2023performance, pu2014effects, ruiz2023degrees, do2023parabolic, xu2017degrees, yuan2021electromagnetic, xu2023exploiting, guo2025impact]:
@@
where $A_{t}$ and $A_{r}$ are the physical areas of Tx and Rx arrays, $D$ is the Tx–Rx separation (treated as a single representative distance), $n$ depends on aperture dimensionality, and $\alpha$ is coefficients determined via orientation of Tx and Rx arrays. More elaborate variants appear in [wang2025analytical]. Our distance-domain result in (38) differs in several key ways:
@@ IV-C2 Comparison with Distance-Domain DoF in the Literature
Let us consider [ding2022degrees], which analyzes two perpendicular ULAs—a geometry closest to our setting. If the Tx array center lies on the Rx-line direction, their closed-form DoF can be written as [ding2022degrees, Eqs. (39),(33)]
@@
Figure 6: Distance-domain DoF vs length of Rx array obtained via approaches: proposed analytical DoF, [ding2022degrees]. and simulation using SVD (with threshold $\epsilon=0.1\sigma_{\max}$). Tx array is ULA with length of $100\lambda$, and $r_{\min}=100\lambda$, while the length of Rx array ($r_{\max}-r_{\min}$) ranges between $10\lambda,2000\lambda]$.
@@
in agreement with simulations (Fig. 6). Thus, when users are spread over a wide range of distances, our proposed DoF expression in (38) captures the correct bounded scaling, while DoF expression in (42) from [ding2022degrees] overestimates growth.
Remark: Unlike many P2P–MIMO DoF’s formulas that implicitly require narrow distance-span assumption ($r_{\min}\!\approx\!r_{\max}$), our distance-domain DoF expression in (38) remains well-posed and can accommodate the large distance spread ($r_{\max}\gg r_{\min}$ or $r_{\max}\rightarrow\infty$). This allows our distance-domain DoF expression to capture the MU-MIMO channel where user’s distance separation can be arbitrarily large.
@@ IV-C3 Non-Uniformity of the Eigenvalue Spectrum
A common (often implicit) assumption in DoF estimates is a nearly flat plateau of dominant eigenvalues followed by rapid decay. This is reasonable when the channel kernel is (approximately) convolutional and the Fourier transform of its kernel is close to an in-band indicator over a passband $\Omega$ (flat in-band) [franceschetti2015landau]:
@@
Implication: Effective-DoF method such as [muharemovic2008antenna, guo2025impact, yuan2021electromagnetic, wang2025analytical, bai2025dynamic]:
@@ V-B Approximation Accuracy for Distance-Domain DoF in the Non-Broadside Case
  • Exact phase & exact amplitude: further incorporates polarization and elementwise distance variation in amplitude [lu2021communicating].
@@
In conclusion, our simulations verify that when $r_{\min}$ exceeds the Fresnel distance (corresponding to a maximum phase error of $\pi/4$), the analytical DoF perfectly matches the numerical results from the exact models. Furthermore, even when $r_{\min}$ falls below this Fresnel distance, the deviation between the analytical prediction and the exact models remains negligible. This confirms that the derived DoF expression is robust against polarization effects and amplitude variations, even when the strict Fresnel conditions are partially relaxed.
@@ VI-B1 Distance-domain DoF under physical length constraint
For this special case, one can see that $\frac{(a+b)}{2}\left(r^{-1}_{\min}-r^{-1}_{\max}\right)$ is approximately the angle subtended by the Rx array seen from the center of the linear Tx sub-array $\mathcal{P}^{(1)}$. Hence, the distance-domain DoF in this special case is analogous to the angular-domain DoF derived for the linear array, as discussed in [poon2005degrees].
As shown by (61), shows that for a fixed sub-array length $L$, increasing the extreme edges (i.e., increasing both $a$ and $b$) expands the distance-domain DoF. Intuitively, increasing the extreme edges expands the angle seen from the center of the sub-array, hence increasing the DoF [poon2005degrees]. This is illustrated in Fig. 14, which suggests that a modular array can exploit additional spatial resources in the distance domain compared to a single-piece array with the same overall physical length.
@@ VI-B2 Spatial DoF under extreme edge constraint
Figure 15: Eigenvalue distributions of LoS channel where the Tx array is either the linear array or the symmetric modular linear array (All outer edges have the same size, and the Rx array satisfies $r_{\min}=200\lambda$ and $r_{\max}=2000\lambda$).
@@ VII Conclusion
In this paper, we investigated the distance-domain DoF by considering a MIMO channel where the source (represents base station) equips with an Tx array with arbitrary shape and the observation region (represents users) is Rx linear array whose elements are aligned at the same angle but lie at different distances from the Tx array. In an ideal scenario where the Tx array is single-piece and the Rx array is on the broadside of the Tx array, we consider both the arrays having continuous aperture with infinite elements and infinitesimal spacing and model the channel as an integral operator with a non-convolution kernel. We further transform it into the integral operator with a Hermitian convolution kernel, which enables eigenvalue analysis via the Fourier transform. This formulation allowed us to derive a closed-form expression for the distance-domain DoF, demonstrating that unlike the angular-domain DoF, the distance-domain DoF does not exhibit linear relationship with aperture’s area and mainly determined by the extreme boundaries of the Tx and Rx arrays rather than its detailed shape.